OpenStudy (anonymous):

I have to answer this question in the next few minutes, anyone have ideas on how? A glider lands 26 miles west and 12 miles south from where it took off. The result of the trip can be described by the vector (-26,-12).

OpenStudy (anonymous):

It is asking what a despription of this vector using distance for magnitude could be

OpenStudy (dumbcow):

magnitude = sqrt((-26)^2 + (-12)^2)

OpenStudy (anonymous):

My only answers are like this: about 29 miles at 25 degrees south of west

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

The magnitude above is the 29. The direction equation for direction is Θ = arctan(y/x) = arctan (12/26) = 25°.